How many representative particles are in 34 g of K 2 O?
1
Expert's answer
2020-12-14T14:50:22-0500

To find the representative particles we need to find no of moles first.

Mole =mass/molar mass

Molar mass of K2O=94

Mole=34/94

Mole=0.362moles

Using the formula

Mole = no of particles/avogadro's constant

No of particles =mole×avogadro's constant

No of particles=0.362×6.02×1023

=2.18×1023